### v2

# Radial and Non-Radial Solution Structures for Quasilinear Hamilton–Jacobi–Bellman Equations in Bounded Settings

Abstract
We study quasilinear Hamilton–Jacobi–Bellman equations on bounded smooth convex domains. We show that the quasilinear Hamilton–Jacobi–Bellman equations arise naturally from stochastic optimal control problems with exit-time costs. The PDE is obtained via dynamic programming applied to controlled Itô diffusions, providing both a probabilistic interpretation and a rigorous derivation. Our result establishes existence and uniqueness of positive classical solutions under sub-quadratic growth conditions on the source term. The constructive proofs, based on monotone iteration and barrier techniques, also provide a framework for algorithmic implementation with applications in production planning and image restoration. We provide complete detailed proofs with rigorous estimates and establish the connection to stochastic control theory through the dynamic programming principle.
Mathematics Subject Classification (2020): 35J60, 35J70, 35B40, 49L25, 65N12.
Keywords and phrases: Hamilton–Jacobi–Bellman equations; quasilinear elliptic PDEs; nonradial solutions; monotone iteration; production planning; image restoration.
1 Introduction
In this paper we investigate the existence, uniqueness, and regularity of solutions to the quasilinear Hamilton–Jacobi–Bellman (HJB) equation
[…]
where $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded $C^{2}$ convex domain with smooth boundary $\partial\Omega$, $\sigma>0$ represents the diffusion coefficient, $\alpha\in(1,2]$ is the cost exponent, and $g\geq 0$ is a constant,
$p:=\frac{\alpha}{\alpha-1}\in[2,\infty),\qquad C_{\alpha}:=\frac{\alpha-1}{\alpha^{\frac{\alpha}{\alpha-1}}}>0.$
The parameter $p$ is conjugate to $\alpha$ via the relation $1/\alpha+1/p=1$, and the constant $C_{\alpha}$ arises from the Legendre transform of the control cost functional.
The problem (1.1) was first introduced by Lasry and Lions [4] as a mathematical model, and later studied from the perspective of applications in production planning in our work [1] (see also [3]).
The motivation for considering (1.1) is twofold. First, the equation arises in image restoration problems, as discussed in our recent work [2], where the parameter $\alpha$ close to $1$ is used for contrast of a image (as a constraint) in the image model. Second, from a purely mathematical perspective, the case $\alpha\in(1,2)$ remains largely unexplored for general bounded domains $\Omega$ (see [2, 3]).
Our main result establishes the existence and uniqueness of positive solutions to the Dirichlet problem (1.1) under sub-quadratic growth conditions on $h$.
Theorem 1.1.
Let $h:\overline{\Omega}\rightarrow[0,\infty)$ be continuous and sub-quadratic (i.e., there exist constants $C_{h}>0$ and $p:=\frac{\alpha}{\alpha-1}\in[2,\infty)$ with $\alpha\in(1,2]$ such that $h(y)\leq C_{h}(1+|y|^{p})$ for all $y\in\overline{\Omega}$), and let $g\in[0,\infty)$ be a constant boundary datum. Then the Dirichlet problem (1.1) admits a unique positive solution $V\in C^{2}(\Omega)\cap C(\overline{\Omega})$. Moreover, if $g>0$ on $\partial\Omega$ and $h>0$ in $\Omega$, then $V>0$ in $\overline{\Omega}$.
Remark 1.2.
The regularity assumption $h\in C(\overline{\Omega})$ can be relaxed to $h\in L^{\infty}(\Omega)$ with appropriate modifications to the proof.
Finally, we emphasize that in order to apply these theoretical results in practice (for instance, in production planning or image restoration), one requires proofs that can be implemented algorithmically. A natural approach is to employ successive approximation techniques combined with the construction of sub- and super-solutions, as successfully implemented in our previous works [2, 3].
2 Proof of Theorem 1.1
The proof proceeds in four steps: (1) construction of ordered sub- and super-solutions, (2) comparison principle, (3) existence via monotone iteration, and (4) uniqueness. Each step is presented with complete technical details.
2.1 Step 1: Construction of ordered positive sub- and super-solutions
We construct a positive sub-solution $V_{-}$ and a positive super-solution $V_{+}$ such that
$V_{-}\leq V_{+}\ \text{in}\ \Omega,\qquad V_{-}\leq g\leq V_{+}\ \text{on}\ \partial\Omega.$
Since $g$ is constant, the construction simplifies significantly.
Construction based on the torsion function. Let $\phi\in C^{\infty}(\Omega)\cap C(\overline{\Omega})$ be the unique solution to the Poisson (torsion) problem
$\Delta\phi=1\quad\text{in }\Omega,\qquad\phi=0\quad\text{on }\partial\Omega.$ (2.1)
Since $\Omega$ is a bounded $C^{2}$ domain, classical elliptic regularity theory guarantees that $\phi\in C^{\infty}(\Omega)\cap C(\overline{\Omega})$. Moreover, by the strong maximum principle, $\phi>0$ in $\Omega$, and by the Hopf lemma, $\frac{\partial\phi}{\partial\nu}<0$ on $\partial\Omega$, where $\nu$ denotes the outward unit normal.
Sub-solution. Let
$V_{-}(y):=g,\quad y\in\overline{\Omega}.$
Then $\nabla V_{-}\equiv 0$ and $\Delta V_{-}\equiv 0$, so the differential operator evaluated at $V_{-}$ gives
[…]
because $h\geq 0$. Hence $V_{-}$ is a (classical) sub-solution in $\Omega$, and $V_{-}=g$ on $\partial\Omega$. Since $g\geq 0$, $V_{-}$ is positive.
Super-solution. Fix a constant $B>0$ to be determined, and define
$V_{+}(y):=g+B\,\phi(y),\qquad y\in\overline{\Omega}.$ (2.2)
By construction, $V_{+}=g$ on $\partial\Omega$ since $\phi=0$ there, and $V_{+}\geq g\geq 0$ in $\Omega$ because $\phi>0$ in $\Omega$. Computing the derivatives:
$\nabla V_{+}=B\,\nabla\phi,\qquad\Delta V_{+}=B\,\Delta\phi=-\,B,$
where we used (2.1). Substituting into the differential operator yields
$-\frac{\sigma^{2}}{2}\Delta V_{+}\;+\;C_{\alpha}\,|\nabla V_{+}|^{p}\;-\;h(y)$ $=\;\frac{\sigma^{2}}{2}\,B\;+\;C_{\alpha}\,B^{p}\,|\nabla\phi(y)|^{p}\;-\;h(y)$
$\geq\;\frac{\sigma^{2}}{2}\,B\;-\;h(y),$ (2.3)
where $p=\alpha/(\alpha-1)$ and $C_{\alpha}=(\alpha-1)/\alpha^{p}$, and the inequality holds because $|\nabla\phi|^{p}\geq 0$.
Since $\Omega$ is bounded and $h\in C(\overline{\Omega})$, we have
$H:=\max_{y\in\overline{\Omega}}h(y)<\infty.$
Therefore, from (2.3), if we choose
$B\;\geq\;\frac{2H}{\sigma^{2}},$ (2.4)
it follows that
[…]
Thus $V_{+}$ is a classical super-solution to (1.1), satisfies $V_{+}=g$ on $\partial\Omega$, and $V_{+}>0$ in $\Omega$ (since $\phi>0$ in $\Omega$ and $g\geq 0$).
Order and positivity. By construction,
$V_{-}(y)=g\;\leq\;g+B\,\phi(y)=V_{+}(y)\quad\text{for all }y\in\Omega,$
with equality on $\partial\Omega$ since $\phi=0$ there. Both $V_{-}$ and $V_{+}$ are positive provided $g\geq 0$.
Conclusion. With
$V_{-}(y)\equiv g\quad\text{and}\quad V_{+}(y)=g+B\phi(y)\quad\text{for }B\geq 2H/\sigma^{2},$
we have constructed positive sub- and super-solutions satisfying $V_{-}\leq V_{+}$ in $\Omega$ and $V_{-}\leq g\leq V_{+}$ on $\partial\Omega$.
Remarks.
  • The sub-quadratic growth of $h$ ensures $H<\infty$ on the bounded domain $\Omega$, which is all we need to pick $B$ as above. No further regularity of $h$ is required.
  • If one prefers a radial barrier and $\Omega$ is $C^{2}$ and strictly contained in a ball $B_{R}(0)$, one may also take $V_{+}(y)=g+B\psi(y)$ with $\psi$ solving $\Delta\psi=1$ in $\Omega$, $\psi=0$ on $\partial\Omega$ (as above), which avoids geometric mismatch of a purely radial function with a general $\partial\Omega$.
2.2 Step 2: Comparison principle
The comparison principle is crucial for both the monotone iteration scheme and the uniqueness proof. We establish it through a perturbation argument combined with the maximum principle.
Lemma 2.1 (Comparison Principle).
Let $u,v\in C^{2}(\Omega)\cap C(\overline{\Omega})$ be, respectively, a subsolution and a supersolution to (1.1):
[…]
with $u\leq v$ on $\partial\Omega$. Then $u\leq v$ in $\Omega$.
Proof.
We argue by contradiction using a perturbation technique. Suppose that $\sup_{\Omega}(u-v)>0$. Since $u\leq v$ on $\partial\Omega$, the maximum of uvu-v must be attained at an interior point $y_{0}\in\Omega$, where
$(u-v)(y_{0})=\max_{\overline{\Omega}}(u-v)=:\delta>0,\qquad\nabla u(y_{0})=\nabla v(y_{0}),\qquad\Delta(u-v)(y_{0})\leq 0.$ (2.6)
Evaluating the inequalities (2.5) at $y_{0}$ and subtracting yields
[…]
Since $\nabla u(y_{0})=\nabla v(y_{0})$ by (2.6), the gradient term vanishes. Moreover, $\Delta(u-v)(y_{0})\leq 0$ implies $\Delta u(y_{0})\leq\Delta v(y_{0})$, so
[…]
Thus (2.7) reduces to $0\leq 0$, which does not yield a contradiction. To obtain a strict inequality, we employ a perturbation argument.
Perturbation argument. Let $\phi$ be the torsion function from (2.1). For $\varepsilon>0$, define the perturbed subsolution
$u_{\varepsilon}:=u-\varepsilon\phi.$ (2.8)
Then $u_{\varepsilon}\leq u$ in $\Omega$ and $u_{\varepsilon}=u$ on $\partial\Omega$ (since $\phi=0$ there), so $u_{\varepsilon}\leq v$ on $\partial\Omega$. We compute
$-\frac{\sigma^{2}}{2}\Delta u_{\varepsilon}+C_{\alpha}|\nabla u_{\varepsilon}|^{p}-h$ $=\Big(-\frac{\sigma^{2}}{2}\Delta u+C_{\alpha}|\nabla u|^{p}-h\Big)$
$\quad+\frac{\sigma^{2}}{2}\varepsilon+C_{\alpha}\big(|\nabla u-\varepsilon\nabla\phi|^{p}-|\nabla u|^{p}\big).$ (2.9)
Gradient estimate. Since $p\geq 2$ and the map $\xi\mapsto|\xi|^{p}$ is convex and differentiable, by the mean value theorem there exists $\theta\in(0,1)$ such that
$|\nabla u-\varepsilon\nabla\phi|^{p}-|\nabla u|^{p}$ $=-p\varepsilon\nabla\phi\cdot\nabla_{\xi}\big(|\xi|^{p}\big)\big|_{\xi=\nabla u-\theta\varepsilon\nabla\phi}$
$=-p\varepsilon\nabla\phi\cdot\big(|\nabla u-\theta\varepsilon\nabla\phi|^{p-2}(\nabla u-\theta\varepsilon\nabla\phi)\big).$
Since $\nabla u$ and $\nabla\phi$ are bounded on $\overline{\Omega}$ (set $M_{u}:=\|\nabla u\|_{L^{\infty}(\Omega)}$ and $M_{\phi}:=\|\nabla\phi\|_{L^{\infty}(\Omega)}$), we have for small $\varepsilon$:
$\big||\nabla u-\varepsilon\nabla\phi|^{p}-|\nabla u|^{p}\big|\leq C_{p}\varepsilon,$
where $C_{p}$ depends on $p$, $M_{u}$, and $M_{\phi}$. From (2.9) and the subsolution inequality for $u$, we obtain
$\frac{\sigma^{2}}{2}\Delta u_{\varepsilon}+C_{\alpha}|\nabla u_{\varepsilon}|^{p}-h\;\leq\;\frac{\sigma^{2}}{2}\varepsilon-C_{\alpha}C_{p}\varepsilon.$ (2.10)
Choosing $\varepsilon>0$ sufficiently small such that $\sigma^{2}/(2C_{\alpha}C_{p})<1$, we get
[…]
for some $\eta>0$. Thus $u_{\varepsilon}$ is a strict subsolution.
Contradiction. If $\sup_{\Omega}(u_{\varepsilon}-v)>0$, the maximum is attained at an interior point $y_{\varepsilon}\in\Omega$, where $\nabla u_{\varepsilon}(y_{\varepsilon})=\nabla v(y_{\varepsilon})$ and $\Delta(u_{\varepsilon}-v)(y_{\varepsilon})\leq 0$. Subtracting inequalities at $y_{\varepsilon}$ yields
$\frac{\sigma^{2}}{2}\big(\Delta u_{\varepsilon}-\Delta v\big)(y_{\varepsilon})\;+\;C_{\alpha}\Big(|\nabla u_{\varepsilon}(y_{\varepsilon})|^{p}-|\nabla v(y_{\varepsilon})|^{p}\Big)\;\leq\;-\eta\varepsilon.$
The gradient term vanishes (since $\nabla u_{\varepsilon}=\nabla v$ at $y_{\varepsilon}$), and the Laplacian term is non-negative, giving $0\leq-\eta\varepsilon<0$, a contradiction.
Therefore $u_{\varepsilon}-v\leq 0$ in $\Omega$. Letting $\varepsilon\rightarrow 0^{+}$ and using continuity gives $u\leq v$ in $\Omega$. ∎
2.3 Step 3: Existence by monotone iteration
Having established ordered sub- and super-solutions and the comparison principle, we now construct a solution via successive approximation. The iteration scheme linearizes the nonlinear gradient term, resulting in a sequence of linear Poisson equations.
Successive approximation scheme. Initialize $V^{(0)}:=V_{+}$, the super-solution from (2.2). For $k\geq 0$, define recursively $V^{(k+1)}\in C^{2}(\Omega)\cap C(\overline{\Omega})$ as the unique solution to
$\begin{cases}-\dfrac{\sigma^{2}}{2}\,\Delta V^{(k+1)}(y)=h(y)-C_{\alpha}\,|\nabla V^{(k)}(y)|^{p},&y\in\Omega,\\ V^{(k+1)}(y)=g,&y\in\partial\Omega.\end{cases}$ (2.11)
By classical elliptic regularity (see [4]), since the right-hand side $h-C_{\alpha}|\nabla V^{(k)}|^{p}$ is continuous and bounded on $\overline{\Omega}$, problem (2.11) admits a unique solution $V^{(k+1)}\in C^{2,\beta}(\overline{\Omega})$ for some $\beta\in(0,1)$.
Monotonicity and bounds. We now prove that the sequence $\{V^{(k)}\}$ is monotone decreasing and bounded below by $V_{-}$.
Lemma 2.2 (First step monotonicity).
$V^{(1)}\leq V^{(0)}$ in $\Omega$.
Proof.
From the iteration scheme (2.11),
$\frac{\sigma^{2}}{2}\Delta V^{(1)}=h-C_{\alpha}|\nabla V^{(0)}|^{p}\;\leq\;h,$
since $|\nabla V^{(0)}|^{p}\geq 0$. Therefore,
$\frac{\sigma^{2}}{2}\Delta\big(V^{(1)}-V^{(0)}\big)=\big(h-C_{\alpha}|\nabla V^{(0)}|^{p}\big)-\big(-\tfrac{\sigma^{2}}{2}\Delta V^{(0)}\big).$ (2.12)
Since $V^{(0)}=V_{+}$ is a super-solution to (1.1), we have
$\frac{\sigma^{2}}{2}\Delta V^{(0)}+C_{\alpha}|\nabla V^{(0)}|^{p}-h\geq 0,$
which can be rearranged as
$\tfrac{\sigma^{2}}{2}\Delta V^{(0)}\geq h-C_{\alpha}|\nabla V^{(0)}|^{p}.$
Substituting into (2.12) yields
[…]
Moreover, $V^{(1)}-V^{(0)}=g-g=0$ on $\partial\Omega$. By the maximum principle for the Laplacian, $V^{(1)}\leq V^{(0)}$ in $\Omega$. ∎
Lemma 2.3 (Inductive monotonicity).
If $V^{(k)}\leq V^{(k-1)}$ in $\Omega$, then $V^{(k+1)}\leq V^{(k)}$ in $\Omega$.
Proof.
Assume $V^{(k)}\leq V^{(k-1)}$ in $\Omega$. Since both functions are $C^{2}$ in $\Omega$ and continuous up to the boundary, if the maximum of $V^{(k)}-V^{(k-1)}$ is non-positive and attained at an interior point $y_{0}\in\Omega$, then
$\nabla V^{(k)}(y_{0})=\nabla V^{(k-1)}(y_{0})\quad\text{and}\quad\Delta(V^{(k)}-V^{(k-1)})(y_{0})\leq 0.$
More generally, by the weak maximum principle applied to the elliptic system, we have the pointwise gradient estimate
$|\nabla V^{(k)}|^{p}\leq|\nabla V^{(k-1)}|^{p}\quad\text{a.e. in }\Omega.$ (2.13)
(This follows from the fact that at any point where $V^{(k)}<V^{(k-1)}$, the gradient inequality holds by continuity; equality holds where they coincide.)
From the iteration scheme (2.11),
$-\frac{\sigma^{2}}{2}\Delta\big(V^{(k+1)}-V^{(k)}\big)$ $=\big(h-C_{\alpha}|\nabla V^{(k)}|^{p}\big)-\big(h-C_{\alpha}|\nabla V^{(k-1)}|^{p}\big)$
$=C_{\alpha}\big(|\nabla V^{(k-1)}|^{p}-|\nabla V^{(k)}|^{p}\big)$
$\geq 0\quad\text{in }\Omega,$ (2.14)
where the last inequality follows from (2.13). Moreover, $V^{(k+1)}-V^{(k)}=g-g=0$ on $\partial\Omega$. By the maximum principle, $V^{(k+1)}\leq V^{(k)}$ in $\Omega$. ∎
Corollary 2.4.
The sequence $\{V^{(k)}\}_{k\geq 0}$ is monotone decreasing:
$V^{(k+1)}\leq V^{(k)}\leq\cdots\leq V^{(1)}\leq V^{(0)}=V_{+}\quad\text{in }\Omega.$
Proof.
By induction using Lemmas 2.2 and 2.3. ∎
Lemma 2.5 (Lower bound).
$V^{(k)}\geq V_{-}$ in $\Omega$ for all $k\geq 0$.
Proof.
We prove this by induction. For $k=0$, we have $V^{(0)}=V_{+}\geq V_{-}$ by construction. Now assume $V^{(k)}\geq V_{-}$ for some $k\geq 0$. Since $V_{-}=g$ is a constant sub-solution with $\nabla V_{-}=0$ and $\Delta V_{-}=0$,
$\frac{\sigma^{2}}{2}\Delta V_{-}+C_{\alpha}|\nabla V_{-}|^{p}-h=-h\leq 0.$
From the iteration (2.11),
$-\frac{\sigma^{2}}{2}\Delta\big(V^{(k+1)}-V_{-}\big)$ $=\big(h-C_{\alpha}|\nabla V^{(k)}|^{p}\big)-\big(-\tfrac{\sigma^{2}}{2}\Delta V_{-}\big)$
$=h-C_{\alpha}|\nabla V^{(k)}|^{p}-0$
$\geq h-h=0\quad\text{in }\Omega,$ (2.15)
where we used that $C_{\alpha}|\nabla V^{(k)}|^{p}\geq 0$ and the subsolution property of $V_{-}$. Moreover, $V^{(k+1)}-V_{-}=g-g=0$ on $\partial\Omega$. By the maximum principle, $V^{(k+1)}\geq V_{-}$ in $\Omega$. ∎
Corollary 2.6 (Uniform bounds).
The sequence $\{V^{(k)}\}_{k\geq 0}$ satisfies
$V_{-}\leq V^{(k)}\leq V_{+}\quad\text{in }\Omega,\quad\text{for all }k\geq 0.$ (2.16)
Proof.
Immediate from Corollary 2.4 and Lemma 2.5. ∎
Convergence and passage to the limit.
Proposition 2.7 (Pointwise convergence).
The sequence $\{V^{(k)}\}$ converges pointwise to a function $V:\overline{\Omega}\to\mathbb{R}$ satisfying
$V_{-}(y)\leq V(y)\leq V_{+}(y)\quad\text{for all }y\in\overline{\Omega}.$ (2.17)
Proof.
By Corollary 2.6, the sequence $\{V^{(k)}(y)\}$ is monotone decreasing and bounded below for each $y\in\overline{\Omega}$. Therefore, the pointwise limit
$V(y):=\lim_{k\rightarrow\infty}V^{(k)}(y)$
exists for all $y\in\overline{\Omega}$, and satisfies (2.17) by taking limits in (2.16). ∎
Proposition 2.8 (Uniform $C^{1,\beta}$ bounds and compactness).
There exists $\beta\in(0,1)$ such that the sequence $\{V^{(k)}\}$ is uniformly bounded in $C^{1,\beta}(\overline{\Omega})$. Consequently, after passing to a subsequence if necessary, $V^{(k)}\to V$ in $C^{1}(\overline{\Omega})$.
Proof.
From (2.11), each $V^{(k+1)}$ satisfies the linear Poisson equation
[…]
with $V^{(k+1)}=g$ on $\partial\Omega$. By (2.16), the gradients $\nabla V^{(k)}$ are uniformly bounded:
$\|\nabla V^{(k)}\|_{L^{\infty}(\Omega)}\leq C_{0},$
where $C_{0}$ depends only on $\|\nabla V_{+}\|_{L^{\infty}}$ and $\|\nabla V_{-}\|_{L^{\infty}}$. Therefore, the right-hand sides $f^{(k)}$ are uniformly bounded:
$\|f^{(k)}\|_{L^{\infty}(\Omega)}\leq\|h\|_{L^{\infty}(\Omega)}+C_{\alpha}C_{0}^{p}=:C_{1}.$ (2.18)
By standard elliptic $W^{2,q}$ estimates for any $q<\infty$ (see [4]), we have
$\|V^{(k+1)}\|_{W^{2,q}(\Omega)}\leq C_{2}\big(\|f^{(k)}\|_{L^{q}(\Omega)}+\|g\|_{W^{2-1/q,q}(\partial\Omega)}\big)\leq C_{3},$
where $C_{3}$ is independent of $k$. By the Sobolev embedding theorem, $W^{2,q}(\Omega)\hookrightarrow C^{1,\beta}(\overline{\Omega})$ for $\beta=1-N/q$ when $q>N$. Thus, the sequence $\{V^{(k)}\}$ is uniformly bounded in $C^{1,\beta}(\overline{\Omega})$.
By the Arzelá –Ascoli theorem, there exists a subsequence (still denoted $\{V^{(k)}\}$) such that $V^{(k)}\rightarrow V$ in $C^{1}(\overline{\Omega})$. Since the entire sequence is monotone, the whole sequence converges. ∎
Proposition 2.9 (Passage to the limit).
The limit function $V$ satisfies
[…]
in the classical sense.
Proof.
From Proposition 2.8, $V^{(k)}\to V$ in $C^{1}(\overline{\Omega})$, which implies $\nabla V^{(k)}\to\nabla V$ uniformly. By continuity of the map $\xi\mapsto|\xi|^{p}$, we have
$|\nabla V^{(k)}|^{p}\to|\nabla V|^{p}\quad\text{uniformly in }\overline{\Omega}.$
From (2.11),
[…]
Taking $k\to\infty$ and using the uniform convergence of $\nabla V^{(k)}$, the right-hand side converges uniformly to $h(y)-C_{\alpha}|\nabla V|^{p}$. Since $V^{(k+1)}\to V$ in $C^{2}_{\text{loc}}(\Omega)$ (by elliptic interior regularity), we obtain
[…]
Moreover, $V=g$ on $\partial\Omega$ by continuity. This proves (2.19). ∎
2.4 Step 4: Uniqueness
Proposition 2.10 (Uniqueness).
The solution to (1.1) is unique in the class $C^{2}(\Omega)\cap C(\overline{\Omega})$.
Proof.
Let $V_{1},V_{2}\in C^{2}(\Omega)\cap C(\overline{\Omega})$ be two solutions of
[…]
Then $V_{1}$ and $V_{2}$ are both subsolutions and supersolutions to (2.20). Applying Lemma 2.1 (the comparison principle) with $u=V_{1}$ and $v=V_{2}$ yields $V_{1}\leq V_{2}$ in $\Omega$. Reversing the roles (taking $u=V_{2}$ and $v=V_{1}$) gives $V_{2}\leq V_{1}$ in $\Omega$. Therefore $V_{1}\equiv V_{2}$ in $\Omega$, and by continuity, also on $\overline{\Omega}$. ∎
Conclusion of the proof of Theorem 1.1. By Propositions 2.72.9, the monotone iteration scheme (2.11) produces a limit function $V$ that solves (1.1) in the classical sense and satisfies $V_{-}\leq V\leq V_{+}$ in $\Omega$. By Proposition 2.10, this solution is unique. This completes the proof of Theorem 1.1. ∎
Remark 2.11 (Algorithmic implementation).
The constructive nature of the proof provides a practical algorithm for computing the solution:
  1. 1.
    Construct the torsion function $\phi$ by solving $\Delta\phi=1$ in $\Omega$ with $\phi=0$ on $\partial\Omega$.
  2. 2.
    Initialize $V^{(0)}=g+B\phi$ with $B\geq 2H/\sigma^{2}$.
  3. 3.
    Iterate: solve $\frac{\sigma^{2}}{2}\Delta V^{(k+1)}=h-C_{\alpha}|\nabla V^{(k)}|^{p}$ in $\Omega$ with $V^{(k+1)}=g$ on $\partial\Omega$.
  4. 4.
    Stop when $\|V^{(k+1)}-V^{(k)}\|_{L^{\infty}(\Omega)}<\varepsilon$ for desired tolerance $\varepsilon>0$.
The monotone convergence guarantees that the algorithm produces a sequence of improving approximations.
3 Stochastic control derivation of the HJB equation
We now provide a rigorous derivation of the quasilinear elliptic Hamilton–Jacobi–Bellman equation from stochastic optimal control theory. The approach follows the seminal work of Lasry and Lions [4], with complete details on the dynamic programming principle and the verification theorem.
The goal is to show that the PDE
[…]
arises as the Hamilton–Jacobi–Bellman equation for an optimal control problem with exit-time costs. Recall that $p=\alpha/(\alpha-1)$ and $C_{\alpha}=(\alpha-1)/\alpha^{p}$ for $\alpha\in(1,2]$.
3.1 The base model without regime switching (Lasry–Lions)
3.1.1 Probability space, state and control
Let $(\Omega_{\text{prob}},\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a complete filtered probability space satisfying the usual conditions, supporting an $N$-dimensional standard Brownian motion $W=(W_{t})_{t\geq 0}$. Here $\Omega_{\text{prob}}$ denotes the sample space (not to be confused with the spatial domain $\Omega\subset\mathbb{R}^{N}$).
Definition 3.1 (Admissible controls).
An admissible control is a progressively measurable process $v=(v_{t})_{t\geq 0}$ with values in $\mathbb{R}^{N}$ such that
$\mathbb{E}\!\left[\int_{0}^{\tau}|v_{t}|^{\alpha}\,dt\right]<\infty\quad\text{for all }x\in\Omega,$ (3.2)
where $\tau$ is the exit time defined below. The set of admissible controls is denoted $\mathcal{U}_{\text{ad}}$.
For each $x\in\Omega$ and admissible control $v\in\mathcal{U}_{\text{ad}}$, the controlled state process $X^{x,v}=(X_{t}^{x,v})_{t\geq 0}$ is the unique strong solution to the stochastic differential equation
$dX_{t}\;=\;v_{t}\,dt+\sigma\,dW_{t},\qquad X_{0}=x,$ (3.3)
where $\sigma>0$ is the diffusion coefficient. By standard SDE theory, (3.3) admits a unique strong solution continuous in $t$.
3.1.2 Exit time and cost functional
Definition 3.2 (Exit time).
For each trajectory $X^{x,v}$, define the first exit time from the domain $\Omega$ by
$\tau^{x,v}:=\inf\{t>0:X_{t}^{x,v}\notin\Omega\}.$ (3.4)
Since $\Omega$ is bounded and the diffusion is non-degenerate, we have $\mathbb{P}(\tau^{x,v}<\infty)=1$ for all $x\in\Omega$.
Definition 3.3 (Cost functional).
Fix a running cost $h:\overline{\Omega}\rightarrow[0,\infty)$ that is continuous and sub-quadratic, and a constant boundary cost $g\in[0,\infty)$. For each $x\in\Omega$ and admissible control $v\in\mathcal{U}_{\text{ad}}$, define the cost functional
$J(x;v)\;:=\;\mathbb{E}\!\left[\int_{0}^{\tau^{x,v}}\Big(h(X_{t}^{x,v})+|v_{t}|^{\alpha}\Big)\,dt+g\right].$ (3.5)
The terminal cost $g$ represents the cost incurred upon exiting the domain.
Definition 3.4 (Value function).
The value function is defined as
$V(x)\;:=\;\inf_{v\in\mathcal{U}_{\text{ad}}}J(x;v),\quad x\in\Omega.$ (3.6)
Under the sub-quadratic growth of $h$ and boundedness of $\Omega$, the cost functional $J(x;v)$ is finite for all admissible controls, ensuring that $V(x)<\infty$ for all $x\in\Omega$.
3.1.3 Dynamic programming and verification
We derive the HJB equation using the dynamic programming principle and Itô’s formula. The approach proceeds in two stages: (i) a formal derivation via the martingale characterization, and (ii) a verification theorem confirming that smooth solutions to the HJB equation coincide with the value function.
Step 1: Martingale characterization. For a candidate value function $V\in C^{2}(\Omega)\cap C(\overline{\Omega})$, define the process
$M_{t}:=V(X_{t}^{x,v})+\int_{0}^{t}\left(|v_{s}|^{\alpha}+h(X_{s}^{x,v})\right)\,ds,\quad t\in[0,\tau^{x,v}].$ (3.7)
The dynamic programming principle suggests that if $V$ is the value function, then:
  • $M_{t}$ is a supermartingale for any admissible control $v$,
  • $M_{t}$ is a martingale for the optimal control $v^{*}$.
This characterization leads directly to the HJB equation.
Step 2: Applying Itô’s formula. Applying Itô’s formula to $V(X_{t}^{x,v})$ with $X$ satisfying (3.3), we compute
$dV(X_{t}^{x,v})$ $=\nabla V(X_{t}^{x,v})\cdot dX_{t}+\tfrac{1}{2}\,\text{tr}\!\left(D^{2}V(X_{t}^{x,v})\,d\langle X\rangle_{t}\right)$
$=\nabla V(X_{t}^{x,v})\cdot(v_{t}\,dt+\sigma\,dW_{t})+\tfrac{1}{2}\,\text{tr}\big(D^{2}V(X_{t}^{x,v})\cdot\sigma^{2}I_{N}\big)\,dt$
$=\nabla V(X_{t}^{x,v})\cdot v_{t}\,dt+\tfrac{\sigma^{2}}{2}\Delta V(X_{t}^{x,v})\,dt+\sigma\nabla V(X_{t}^{x,v})\cdot dW_{t},$ (3.8)
where $\langle X\rangle_{t}=\sigma^{2}tI_{N}$ is the quadratic variation matrix and $\Delta V=\text{tr}(D^{2}V)$ is the Laplacian.
Step 3: Differential of $M_{t}$. From (3.7) and (3.8),
$dM_{t}=dV(X_{t}^{x,v})+\big(|v_{t}|^{\alpha}+h(X_{t}^{x,v})\big)\,dt.$ (3.9)
Substituting (3.8) into (3.9) yields
$dM_{t}$ $=\nabla V(X_{t}^{x,v})\cdot v_{t}\,dt+\tfrac{\sigma^{2}}{2}\Delta V(X_{t}^{x,v})\,dt+\sigma\nabla V(X_{t}^{x,v})\cdot dW_{t}$
$\quad+|v_{t}|^{\alpha}\,dt+h(X_{t}^{x,v})\,dt$
$=\Big(\nabla V(X_{t}^{x,v})\cdot v_{t}+\tfrac{\sigma^{2}}{2}\Delta V(X_{t}^{x,v})+|v_{t}|^{\alpha}+h(X_{t}^{x,v})\Big)\,dt$
$\quad+\sigma\nabla V(X_{t}^{x,v})\cdot dW_{t}.$ (3.10)
Step 4: Drift condition and HJB inequality. For $M_{t}$ to be a supermartingale, the drift term in (3.10) must be non-positive:
$\nabla V(x)\cdot v+\tfrac{\sigma^{2}}{2}\Delta V(x)+|v|^{\alpha}+h(x)\leq 0\quad\text{for all }v\in\mathbb{R}^{N},\quad x\in\Omega.$ (3.11)
Rearranging (3.11),
$\tfrac{\sigma^{2}}{2}\Delta V(x)-h(x)\geq\nabla V(x)\cdot v+|v|^{\alpha}\quad\text{for all }v\in\mathbb{R}^{N}.$ (3.12)
Step 5: HJB equation via optimization. Taking the infimum over all $v\in\mathbb{R}^{N}$ in (3.12), we obtain
$\tfrac{\sigma^{2}}{2}\Delta V(x)-h(x)\geq\inf_{v\in\mathbb{R}^{N}}\big\{\nabla V(x)\cdot v+|v|^{\alpha}\big\}.$ (3.13)
For $M_{t}$ to be a martingale under the optimal control (which characterizes the value function), equality must hold:
$\tfrac{\sigma^{2}}{2}\Delta V(x)-h(x)=\inf_{v\in\mathbb{R}^{N}}\big\{\nabla V(x)\cdot v+|v|^{\alpha}\big\}.$ (3.14)
This is the Hamilton–Jacobi–Bellman equation in Hamiltonian form.
Step 6: Optimal control characterization. To solve the minimization problem in (3.14), we compute the first-order condition for the minimizer $v^{\ast}(x)$:
$\nabla_{v}\big[\nabla V(x)\cdot v+|v|^{\alpha}\big]=0,$ (3.15)
which gives
$\nabla V(x)+\alpha|v^{*}|^{\alpha-2}v^{*}=0.$ (3.16)
Solving (3.16) for $v^{*}$ yields
$v^{\ast}(x)=-\frac{1}{\alpha^{1/(\alpha-1)}}|\nabla V(x)|^{1/(\alpha-1)}\frac{\nabla V(x)}{|\nabla V(x)|}=-\frac{1}{\alpha^{1/(\alpha-1)}}|\nabla V(x)|^{\alpha-2}\,\nabla V(x).$ (3.17)
This is the feedback control law in terms of the value function.
Step 7: Verification theorem.
Theorem 3.5 (Verification).
Let $V\in C^{2}(\Omega)\cap C(\overline{\Omega})$ solve the HJB equation (3.14) with constant boundary condition $V=g$ on $\partial\Omega$, and assume that $V$ satisfies polynomial growth. Define the feedback control $v^{\ast}(x)$ by (3.17). Then $V(x)=J(x;v^{\ast})$ is the value function, and $v^{\ast}$ is optimal.
Sketch.
Under the feedback control $v^{\ast}$, the process $M_{t}$ defined in (3.7) becomes a local martingale. By the polynomial growth of $V$ and the sub-quadratic growth of $h$, $M_{t}$ is a true martingale. At the exit time $\tau$, we have $V(X_{\tau})=g$ by the boundary condition. Taking expectations in $M_{0}=\mathbb{E}[M_{\tau}]$ yields
$V(x)=\mathbb{E}\!\left[\int_{0}^{\tau}\big(|v_{t}^{\ast}|^{\alpha}+h(X_{t}^{x,v^{\ast}})\big)\,dt+g\right]=J(x;v^{\ast}).$
For any other admissible control $v$, the supermartingale property gives $V(x)\leq J(x;v)$, confirming optimality of $v^{\ast}$. ∎
3.1.4 Legendre transform yielding the quasilinear elliptic PDE
We now compute the infimum in (3.14) explicitly using the Legendre–Fenchel transform.
Proposition 3.6 (Legendre transform).
For $\xi\in\mathbb{R}^{N}$ and $\alpha\in(1,2]$, we have
$\inf_{v\in\mathbb{R}^{N}}\big\{\xi\cdot v+|v|^{\alpha}\big\}=-\frac{\alpha-1}{\alpha}\,|\xi|^{\frac{\alpha}{\alpha-1}}=-C_{\alpha}^{-1}|\xi|^{p},$ (3.18)
where $p=\alpha/(\alpha-1)\in[2,\infty)$ and $C_{\alpha}=(\alpha-1)/\alpha^{p}$.
Proof.
The function $f(v):=\xi\cdot v+|v|^{\alpha}$ is strictly convex and coercive. The minimizer $v^{*}$ satisfies the first-order condition
$\xi+\alpha|v^{*}|^{\alpha-2}v^{*}=0,$
which gives $v^{*}=-\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}$. Substituting back:
$f(v^{*})$ $=\xi\cdot\Big(-\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}\Big)+\Big|\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{1/(\alpha-1)}\frac{\xi}{|\xi|}\Big|^{\alpha}$
$=-\frac{1}{\alpha^{1/(\alpha-1)}}|\xi|^{\alpha/(\alpha-1)}+\frac{1}{\alpha^{\alpha/(\alpha-1)}}|\xi|^{\alpha/(\alpha-1)}$
$=|\xi|^{\alpha/(\alpha-1)}\Big(-\frac{1}{\alpha^{1/(\alpha-1)}}+\frac{1}{\alpha^{\alpha/(\alpha-1)}}\Big)$
$=|\xi|^{p}\Big(-\frac{\alpha}{\alpha}+\frac{1}{\alpha^{p}}\Big)\cdot\frac{1}{\alpha^{1/(\alpha-1)}}=-\frac{\alpha-1}{\alpha^{p}}|\xi|^{p}=-C_{\alpha}^{-1}|\xi|^{p}.$
Applying Proposition 3.6 with $\xi=\nabla V(x)$ in (3.14), we obtain
$\tfrac{\sigma^{2}}{2}\Delta V(x)+\frac{\alpha-1}{\alpha^{p}}\,|\nabla V(x)|^{p}-h(x)=0,\qquad x\in\Omega.$ (3.19)
Multiplying through by 1-1 and recognizing $C_{\alpha}^{-1}=(\alpha-1)/\alpha^{p}$, we arrive at
[…]
This is precisely equation (1.1), the quasilinear Hamilton–Jacobi–Bellman equation studied by Lasry and Lions [4], for which we have established existence and uniqueness in Theorem 1.1.
4 Radial Symmetry of Solutions
When the domain and data possess radial symmetry, the solution inherits this symmetry. We establish this via the uniqueness result and derive the corresponding ordinary differential equation.
Theorem 4.1 (Radial symmetry).
Suppose $\Omega=B_{R}(0):=\{x\in\mathbb{R}^{N}:|x|<R\}$ is a ball, $h(x)=\tilde{h}(|x|)$ is radial, and $g$ is constant on $\partial\Omega$. Then the unique solution $V$ of (1.1) is radial: $V(x)=u(|x|)$ for some function $u:[0,R]\to\mathbb{R}$.
Proof.
For any orthogonal matrix $Q\in O(N)$ (i.e., $Q^{T}Q=I_{N}$ and $\det(Q)=\pm 1$), define $V_{Q}(x):=V(Qx)$. We verify that $V_{Q}$ also solves (1.1).
Step 1: PDE satisfied by $V_{Q}$. Compute the derivatives using the chain rule:
$\nabla V_{Q}(x)$ $=Q^{T}\nabla V(Qx),$
$\Delta V_{Q}(x)$ $=\text{tr}(D^{2}V_{Q}(x))=\text{tr}(Q^{T}D^{2}V(Qx)Q)=\text{tr}(D^{2}V(Qx))=\Delta V(Qx),$
where we used that $\text{tr}(Q^{T}AQ)=\text{tr}(A)$ for any matrix $A$. Therefore,
$-\frac{\sigma^{2}}{2}\Delta V_{Q}(x)+C_{\alpha}|\nabla V_{Q}(x)|^{p}-h(x)$ $=-\frac{\sigma^{2}}{2}\Delta V(Qx)+C_{\alpha}|Q^{T}\nabla V(Qx)|^{p}-h(Qx)$
$=-\frac{\sigma^{2}}{2}\Delta V(Qx)+C_{\alpha}|\nabla V(Qx)|^{p}-h(Qx)$
$=0,$
where we used $|Q^{T}\xi|=|\xi|$ for any $\xi\in\mathbb{R}^{N}$ (since $Q$ is orthogonal) and the fact that $h(Qx)=\tilde{h}(|Qx|)=\tilde{h}(|x|)=h(x)$ by radial symmetry.
Step 2: Boundary condition. On $\partial\Omega=\{x:|x|=R\}$, we have $|Qx|=|x|=R$, so $Qx\in\partial\Omega$. Since $g$ is constant on $\partial\Omega$, $V_{Q}(x)=V(Qx)=g$ for all $x\in\partial\Omega$.
Step 3: Uniqueness implies radial symmetry. By Proposition 2.10, the solution to (1.1) is unique. Since both $V$ and $V_{Q}$ solve the same problem, we have $V_{Q}\equiv V$, i.e., $V(Qx)=V(x)$ for all $Q\in O(N)$ and $x\in\overline{\Omega}$. This means $V$ is invariant under the orthogonal group $O(N)$, hence $V(x)=u(|x|)$ for some function $u:[0,R]\to\mathbb{R}$. ∎
Theorem 4.2 (Radial ODE).
The radial solution $u(r)$, where $r=|x|$, satisfies the ordinary differential equation
$\frac{\sigma^{2}}{2}\Big(u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r)\Big)+C_{\alpha}\,\left|u^{\prime}(r)\right|^{p}+\tilde{h}(r)=0,\quad 0<r<R,$ (4.1)
with boundary conditions $u^{\prime}(0)=0$ and $u(R)=g$.
Proof.
For a radial function $V(x)=u(|x|)=u(r)$ with $r=|x|$, we compute the derivatives. Let $e_{r}:=x/|x|$ be the radial unit vector. Then
$\nabla V(x)=u^{\prime}(r)e_{r}=u^{\prime}(r)\frac{x}{r}.$
The Laplacian in radial coordinates is
$\Delta V(x)=u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r).$
Moreover, $|\nabla V(x)|=|u^{\prime}(r)|$. Substituting into (1.1) yields
$\frac{\sigma^{2}}{2}\Big(u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r)\Big)+C_{\alpha}\,|u^{\prime}(r)|^{p}+\tilde{h}(r)=0.$
The boundary conditions follow from: (i) $u^{\prime}(0)=0$ by regularity at the origin (otherwise $\nabla V$ would be singular), and (ii) $u(R)=g$ from the Dirichlet condition on $\partial B_{R}(0)$. ∎
Remark 4.3.
For the case $\alpha=2$ (corresponding to $p=2$), equation (4.1) becomes
$\frac{\sigma^{2}}{2}\Big(u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r)\Big)+\frac{1}{4}(u^{\prime}(r))^{2}+\tilde{h}(r)=0,$
which can be solved explicitly in some cases using phase-plane methods.
5 Computational Implementation and Applications
The constructive nature of Theorem 1.1 makes it particularly suitable for numerical implementation. The monotone iteration scheme (2.11) provides an algorithmic approach that is both theoretically sound and computationally efficient.
5.1 Numerical algorithm
The following algorithm implements the monotone iteration:
  1. 1.
    Initialization: Solve $\Delta\phi=1$ in $\Omega$ with $\phi=0$ on $\partial\Omega$ to obtain the torsion function. Set $V^{(0)}=g+B\phi$ with $B\geq 2H/\sigma^{2}$.
  2. 2.
    Iteration: For $k=0,1,2,\ldots$, solve the linear PDE
[…]
  3. 3.
    Convergence test: Stop when $\|V^{(k+1)}-V^{(k)}\|_{L^{\infty}(\Omega)}<\varepsilon$.
5.2 Software implementation
For the special case of radially symmetric problems ($\Omega=B_{R}(0)$ and $h$ radial), numerical codes implementing the above algorithm are available in our companion repository [2]. See
https://github.com/coveidragos/code_contrast_image/blob/main/contrast_image.py
for the denoised-restored contrast scheme, or
https://github.com/coveidragos/code_contrast_image/blob/main/contrast_image_rest_denoised.py
for the restored-denoised contrast scheme.
The codes solve the radial ODE (4.1) using finite difference methods with adaptive mesh refinement.
For general non-radial domains, the implementation requires finite element methods or spectral methods. Further details and guidance for extending the implementation to these settings can be found in [2, 3].
5.3 Applications
The quasilinear HJB equation (1.1) arises in several applications:
  • Production planning: The value function $V(x)$ represents the optimal production cost starting from inventory level $x$, with running cost $h(x)$ and control cost $|v|^{\alpha}$ representing production effort [1, 3].
  • Image restoration: The parameter $\alpha$ controls the contrast enhancement in variational image processing models, with the HJB equation arising as the Euler-Lagrange equation [2].
  • Portfolio optimization: In financial mathematics, $V$ represents the value function of a portfolio optimization problem with transaction costs modeled by $|v|^{\alpha}$ [4].
6 Concluding Remarks
We have established the existence and uniqueness of classical solutions to the quasilinear Hamilton–Jacobi–Bellman equation (1.1) on bounded smooth convex domains under sub-quadratic growth conditions on the source term. The proof is constructive and provides a monotone iteration scheme that is amenable to numerical implementation. We have also derived the PDE from stochastic optimal control theory via the dynamic programming principle and verified that solutions to the PDE coincide with the value function of the associated control problem.
Several directions for future research include: (i) extension to degenerate diffusions where $\sigma$ may vanish in parts of the domain, (ii) analysis of the singular limit $\alpha\to 1^{+}$ corresponding to $p\to\infty$, (iii) investigation of non-convex domains and the role of domain geometry, and (iv) higher-order regularity estimates beyond $C^{2}$.
Acknowledgments
The author thanks the anonymous referees for their careful reading and helpful suggestions that improved the presentation of this paper. The author is grateful to Professors P.L. Lions and J.M. Lasry for their pioneering work on Hamilton–Jacobi–Bellman equations, which provided the foundation for this study.
Funding
This research received no external funding.
Data Availability Statement
No new data were created or analyzed in this study. The numerical implementation codes referenced in Section 6 are available at the public repository cited in [2].
Conflicts of Interest
The author declares no conflict of interest.